# Generalizing Deepfake Video Detection with Plug-and-Play: Video-Level Blending and Spatiotemporal Adapter Tuning

### v1

@@ Abstract
Three key challenges hinder the development of current deepfake video detection: (1) Temporal features can be complex and diverse: how can we identify general temporal artifacts to enhance model generalization? (2) Spatiotemporal models often lean heavily on one type of artifact and ignore the other: how can we ensure balanced learning from both? (3) Videos are naturally resource-intensive: how can we tackle efficiency without compromising accuracy?
This paper attempts to tackle the three challenges jointly. First, inspired by the notable generality of using image-level blending data for image forgery detection, we investigate whether and how video-level blending can be effective in video. We then perform a thorough analysis and identify a previously underexplored temporal forgery artifact: Facial Feature Drift (FFD), which commonly exists across different forgeries. To reproduce FFD, we then propose a novel Video-level Blending data (VB), where VB is implemented by blending the original image and its warped version frame-by-frame, serving as a hard negative sample to mine more general artifacts. Second, we carefully design a lightweight Spatiotemporal Adapter (StA) to equip a pretrained image model (both ViTs and CNNs) with the ability to capture both spatial and temporal features jointly and efficiently. StA is designed with two-stream 3D-Conv with varying kernel sizes, allowing it to process spatial and temporal features separately. Extensive experiments validate the effectiveness of the proposed methods; and show our approach can generalize well to previously unseen forgery videos, even the just-released (in 2024) SoTAs. We release our code and pretrained weights at https://github.com/YZY-stack/StA4Deepfake.
@@ 1 Introduction
Current deepfake [^1] The term “deepfake” hereafter mainly refers to face forgery, typically referring to face-swapping. The entire (or natural) image synthesis is not within our scope. technology enables effortless manipulation of facial identities in video content, bringing many impressive applications in fields such as movie-making and entertainment. However, this technology can also be misused for malicious purposes, including violating personal privacy, spreading misinformation, and eroding trust in digital media. Hence, there is a pressing need to establish a reliable system for detecting deepfake videos.
@@
Most prior deepfake detectors (Zhou et al. 2017; Li and Lyu 2018; Rossler et al. 2019; Yang, Li, and Lyu 2019) perform well within the same dataset. However, they often fail to generalize well in cross-dataset scenarios where training and testing data distributions differ. Training with synthetic (blending) images appears to be one of the most effective solutions to this problem, which encourages detectors to learn generic representations ($e.g.$, blending boundaries and inconsistencies between inner and outer faces) for deepfake detection (Li et al. 2020a; Shiohara and Yamasaki 2022). Given the notable success of image-level blending, we ask: what about video-level blending? Can this method also be applied to learn general forgery artifacts in the video domain? (Research-Question-1).
@@
To this end, we perform a comprehensive examination of current realistic forgery videos and discover an inconsistency in the facial organs between forgery frames (as illustrated in Fig. 1). We term this phenomenon as the Facial Feature Drift (FFD). As illustrated in Fig. 1, even when two consecutive frames appear relatively static (possibly without other unrelated effects), subtle unnatural drift and flickering can be observed within facial features ($e.g.$, eyes, nose, $etc$). This observation provides reasonable evidence that unnatural forgery artifacts are indeed present in the temporal domain of manipulated videos.
@@
Intuitively, the root cause of FFD can be attributed to the frame-by-frame face-swapping operation employed by existing deepfake algorithms. To understand the underlying reasons, we argue that the primary factor is the randomness during the generation process of fake faces. Deepfake algorithms typically create a fake frame by using DNNs to generate a counterfeit inner face. Subsequently, they perform a blending operation to integrate this generated face back into the original background. Given this context, it is important to note that the inner face between different frames within a single generated video may exhibit variations, as the DNN is not a deterministic function. Consequently, this leads to inconsistencies in the generation process of fake faces.
@@
To effectively make the detection model capture FFD in the given deepfake, we introduce a novel Video-level Blending data called VB to reproduce FFD. The overall creation pipeline of VB is illustrated in Fig. 2. Specifically, we generate VB by first applying warping operations separately to facial organs and then performing image-level self-blending to generate one synthesis image. The process is repeatedly conducted for all frames within one clip to create a set of synthesis images (video-level pseudo-fakes). VB can serve as the hard negative samples for mining more general discriminative features for detection. But why do we specifically focus on facial features? This is mainly because regions, such as skin, are low-level areas that can be sensitive to random perturbations ($e.g.$, compression), while facial features typically possess a more robust high-level layout. Also, we can observe more obvious temporal inconsistency within the facial organs in Fig. 1.
@@
Additionally, as suggested by (Zheng et al. 2021; Wang et al. 2023), naively training a temporal detector may result in the model only focusing on easier types of artifacts and overlooking more challenging ones (e.g., learning spatial only). So how do we jointly learn temporal and spatial features without heavily relying on one? (Research-Question-2). Also, as videos are naturally more complex than images with multiple frames to be computed, how do we address the efficiency issue? (Research-Question-3). Instead of designing a new video architecture from scratch, we propose an alternative approach: to “convert" a pretrained image model into a video model, enabling the model to learn additional temporal information. Specifically, we meticulously design a lightweight, plug-and-play Spatiotemporal Adapter (StA) that equips an image model ($e.g.$, ViT or CNN) with spatiotemporal reasoning capabilities. Throughout the whole training process, only the added adapters are updated, which significantly reduces the training cost and allows us to leverage the powerful generalization capabilities of SoTA image models (e.g., CLIP).
@@ 2.1 Deepfake Image Detection
Conventional image detectors (Nguyen, Yamagishi, and Echizen 2019; Rossler et al. 2019) typically focus on developing optimal CNN architectures ($e.g.$, MesoNet, CapsuleNet). However, these methods often overlook the details present in the frequency domain of fake images, such as compression artifacts. To this end, several works (Qian et al. 2020; Gu et al. 2022a) utilize frequency information to improve the performance of detectors. Other notable directions are focusing on some specific representations, i.e., 3D decomposition (Zhu et al. 2021), disentanglement (Liang, Shi, and Deng 2022; Yan et al. 2023b), erasing technology (Wang and Deng 2021), theoretic information (Sun et al. 2022a), debiased learning (Cheng et al. 2024), facial landmark (Yan et al. 2022), and parameter-efficient tuning (Kong, Li, and Wang 2023; Kong et al. 2024b, a). Recently, several works (Jia et al. 2024; Shi et al. 2024) attempt to detect deepfakes with training-free pipelines. These works mostly utilize designed algorithms for detection from the view of representation learning. Another important direction of image forgery detection is data synthesis (image blending). We will discuss it separately in another subsection.
@@ 2.2 Deepfake Video Detection
There are parts of works (Zhou et al. 2017) that focus on temporal artifacts. An early work (Haliassos et al. 2021) (in 2021) proposes to leverage the temporal inconsistency of the mouth movement. This work utilizes a pre-trained model to learn the natural lip movement as prior on a lip-reading dataset. FTCN (Zheng et al. 2021) explores directly training a fully temporal 3D ConvNets with an attached temporal Transformer. However, detecting without spatial information may harm the generalization capability. STIL (Gu et al. 2021) considers both spatial and temporal inconsistency and designs a spatiotemporal inconsistency Learning framework for deepfake video detection. RealForensics (Haliassos et al. 2022) introduces audio information and leverages self-supervised learning for representation learning. Very recently, wang et al.(Wang et al. 2023) propose a new learning method that alternatively trains and freezes the spatial and temporal convolutional layers in a 3D CNN ($e.g.$, I3D R50). The result shows the importance of leveraging both spatial and temporal features for detecting deepfakes.
@@ 2.3 Deepfake Detectors Based on Data Synthesis
Image-level Data Synthesis. One effective approach in deepfake detection is to use synthetic data (or blending data/images) for training. This strategy encourages models to learn generic representations for detecting deepfakes, such as blending boundaries and inconsistency between the inner and outer faces (Shiohara and Yamasaki 2022). For instance, in the early stages, FWA (Li and Lyu 2018) employs a self-blending strategy by applying image transformations (e.g., down-sampling) to the facial region and then warping it back into the original image. This process is designed to learn the wrapping artifacts during the deepfake generation process. Another noteworthy contribution is Face X-ray (Li et al. 2020a), which explicitly encourages detectors to learn the blending boundaries of fake images. Furthermore, SLADD (Chen et al. 2022) introduces an adversarial method to dynamically generate the most challenging blending choices for synthesizing data. Rather than swapping faces between two different identities, a recent art, SBI (Shiohara and Yamasaki 2022), proposes to swap with the same person’s identity to reach a high-realistic face-swapping. Given the notable generalization performance for detecting image forgery by using synthetic data, it remains an open question whether and how video-level synthetic data can be effective.
Video-level Data Synthesis. Compared to image-level data synthesis (blending), very little works pay attention to designing a suitable video-level blending method. To the best of our knowledge, we identify one notable work (Wang et al. 2023) that attempts to design video-level data synthesis by using techniques such as dropping/repeating one frame within one clip. However, it is possible that these common operations might not explicitly represent the general forgery pattern. Recall that in the image domain, we use blending technology to synthesize the forgery artifact since the blending artifact commonly exists across different forgeries, generally. So there are two main challenges: (1) What kinds of features generally exist in the temporal domain? (2) How to use blending to simulate it like the image data synthesis? We will show our video-level method in the following section.
@@ 3.3 Spatiotemporal Adapters (StA)
A video can be naturally represented as a stack of image frames. One challenging problem is to capture both spatial and temporal forgery artifacts without heavily relying on one type of artifact. Previous works generally follow two mainstream frameworks: (1) Utilizing 3D CNN ($e.g.$, I3D R50) for capturing spatiotemporal information from forgery videos (Zheng et al. 2021; Wang et al. 2023); (2) Utilizing two-stream architecture to learn the spatial and temporal features in two separate branches (Masi et al. 2020; Zhou et al. 2017). These two paradigms typically involve fully fine-tuning for both spatial and temporal modules. However, training all layers of both spatial and temporal modules will inevitably introduce a large number of learnable parameters, making the optimization and training procedure challenging. On the other hand, the image model (especially ViT-based image SoTA) is developing rapidly. It is well-known that these image SoTAs have achieved notable performance in several downstream computer vision tasks ($e.g.$, multi-modality (Ganz et al. 2024)). Therefore, can we leverage the power of these SoTA image models to effectively and efficiently finetune the face forgery detection task? The key challenge is to perform the temporal forgery feature modeling in a given image model.
@@
Table 1: Protocol-1: Cross-dataset evaluations with image detectors. All detectors are trained on FF++_c23 (Rossler et al. 2019) and evaluated on other datasets. The best results are highlighted in bold and the second is underlined. * indicates the results are directly cited from their original papers, otherwise the results are reproduced in our benchmark.
@@
F3Net (Qian et al. 2020) ECCV 2020 0.789 0.844 0.718 0.749 0.730 0.728 0.649 0.743
SPSL (Liu et al. 2021) CVPR 2021 0.799 0.871 0.724 0.770 0.723 0.702 0.794 0.769
SRM (Luo et al. 2021) CVPR 2021 0.840 0.885 0.695 0.728 0.722 0.702 0.794 0.767
CORE (Ni et al. 2022) CVPRW 2022 0.809 0.882 0.721 0.720 0.765 0.724 0.710 0.762
RECCE (Cao et al. 2022) CVPR 2022 0.823 0.891 0.696 0.734 0.784 0.756 0.711 0.779
DCL (Sun et al. 2022b) AAAI 2022 0.882 0.921 0.750 0.769 0.975 0.776 0.863 0.851
SLADD* (Chen et al. 2022) CVPR 2022 0.797 - 0.772 - - - - -
SBI (Shiohara and Yamasaki 2022) CVPR 2022 0.886 0.827 0.717 0.848 0.899 0.703 0.866 0.821
UIA-ViT* (Zhuang et al. 2022) ECCV 2022 0.889 - - 0.795 - - - -
UCF (Yan et al. 2023b) ICCV 2023 0.837 0.867 0.742 0.770 0.808 0.774 0.697 0.785
SeeABLE* (Yan et al. 2023b) ICCV 2023 0.873 - 0.759 0.863 - - - -
IID* (Huang et al. 2023) CVPR 2023 0.838 0.939 - 0.812 - - - -
CFM (Luo et al. 2023) TIFS 2023 0.897 0.952 0.706 0.802 0.976 0.823 0.831 0.856
LSDA (Yan et al. 2023a) CVPR 2024 0.898 0.956 0.735 0.812 0.892 0.756 0.879 0.847
@@
Table 2: Protocol-2: Cross-manipulation evaluations within FF++ domain. We use the just-released latest deepfake dataset DF40 (Yan et al. 2024) and select eight typical face-swapping methods generated on the FF++ domain, allowing us to perform the cross-manipulation evaluation (domain unchanged).
@@
RECCE (Cao et al. 2022) 0.898 0.832 0.925 0.683 0.848 0.949 0.848 0.768 0.844
SBI (Shiohara and Yamasaki 2022) 0.724 0.891 0.952 0.750 0.594 0.803 0.712 0.701 0.766
UCF (Yan et al. 2023b) 0.831 0.827 0.950 0.731 0.862 0.937 0.809 0.647 0.824
AltFreezing (Wang et al. 2023) 0.947 0.951 0.851 0.605 0.836 0.958 0.927 0.797 0.859
@@
LipForensics (Haliassos et al. 2021) CVPR 2021 0.824 0.735 0.780
FTCN (Zheng et al. 2021) ICCV 2021 0.869 0.740 0.805
HCIL (Gu et al. 2022b) ECCV 2022 0.790 0.692 0.741
RealForensics (Haliassos et al. 2022) CVPR 2022 0.857 0.759 0.808
LTTD (Guan et al. 2022) NeurIPS 2022 0.893 0.804 0.849
AltFreezing (Wang et al. 2023) CVPR 2023 0.895 - -
TALL-Swin (Xu et al. 2023) ICCV 2023 0.908 0.768 0.838
StyleDFD (Choi et al. 2024) CVPR 2024 0.890 - -
NACO (Zhang et al. 2024) ECCV 2024 0.895 0.767 0.831
@@
4 Results
4.1 Setup
Dataset. To evaluate the generalization ability of our proposed framework, we conduct experiments on eight commonly used datasets: FaceForensics++ (FF++) (Rossler et al. 2019), Deepfake Detection Challenge (DFDC) (Dolhansky et al. 2020), the preview version of DFDC (DFDCP) (Dolhansky et al. 2019), DeepfakeDetection (DFD) (Deepfakedetection. 2021), Celeb-DF-v2 (CDF-v2) (Li et al. 2020b), DeeperForensics (DFo) (Jiang et al. 2020), WildDeepfake (WDF) (Zi et al. 2020), and FFIW (Zhou et al. 2021). We also evaluate the models on the just-released (in 2024) deepfake dataset DF40 (Yan et al. 2024), which contains 40 distinct forgery methods. We select eight face-swapping methods generated from the FF++ domain for our evaluation. In line with the standard deepfake benchmark (Yan et al. 2023c), we utilize the c23 version of FF++ for training and other datasets for testing.
Implementation Details. We utilize ViT-l14 (Dosovitskiy et al. 2020) as the backbone for extracting forgery features. By default, the model parameters are initialized through pre-training on CLIP (Radford et al. 2021). We also investigate other architectures in Tab. 4. In our experiments, we uniformly sample 32 frames from each video during training and randomly select 8 consecutive frames as a clip for each iteration. All preprocessing and training codebases are aligned to the deepfake benchmark (Yan et al. 2023c) for a fair comparison. We also employ several widely used data augmentations, such as CutOut, and ImageCompression.
Evaluation Protocols. We adopt three standard protocols to evaluate our models: Protocol-1: cross-dataset evaluation with image methods, Protocol-2: cross-manipulation evaluation on the FF++ data domain, and Protocol-3: cross-dataset evaluation with video methods.
4.2 Protocol-1: Cross-dataset Evaluation with image detectors
In Tab. 1, we compare our method with several latest and recent SoTA detectors. Most detectors (wo * marker) are reproduced in our benchmark under the same settings for a fair comparison. We see that our method consistently outperforms other models across all tested scenarios. Note that methods with ‘*’ do not have the complete results of all the testing datasets we used, so we leave ‘-’ for these results.
4.3 Protocol-2: Cross-manipulation evaluation
Forgery methods in the real world can be highly diverse. Evaluating our model’s performance on cross-manipulation tasks allows us to determine whether it can handle previously unseen fake types. Instead of training and testing on the older FF++ dataset with four fake methods, we use the recently released DF40 dataset (Yan et al. 2024) for evaluation. For comparison, we select four state-of-the-art detectors, including both image and video models. Our method generally outperforms other models on average, particularly on e4s, Inswap, and SimSwap methods (see Tab. 2). This demonstrates that our method learns more general features for detection, even for the latest techniques.
Figure 4: Robustness to Unseen Perturbations: We present video-level AUC for five distinct degradation levels across five types of perturbations in (Jiang et al. 2020).
@@ 4.3 Protocol-2: Cross-manipulation evaluation
I3D-RGB (Haliassos et al. 2021) 25M 0.625 0.655 0.640
3D R50 (Haliassos et al. 2021) 46M 0.824 0.735 0.780
VTN (Gu et al. 2021) 46M 0.734 0.691 0.713
VidTR (Zheng et al. 2021) 93M 0.869 0.740 0.805
ViViT (Wang et al. 2023) 310M 0.895 - -
ISTVT (Tong et al. 2022) - 0.606 0.551 0.579
ResNet-34 (He et al. 2016) 21M $\times$ 0.742 0.626 0.684
@@
4.4 Protocol-3: Cross-dataset evaluations with video detectors
Forgery methods in the real world can be highly diverse. Evaluating our model’s performance on cross-manipulation tasks allows us to determine whether it can handle previously unseen fake types. Instead of training and testing on the older FF++ dataset with four fake methods, we use the recently released DF40 dataset (Yan et al. 2024) for evaluation. For comparison, we select four state-of-the-art detectors, including both image and video models. Our method generally outperforms other models on average, particularly on e4s, Inswap, and SimSwap methods. This demonstrates that our method learns more general features for detection.
4.5 Robustness Evaluation
To evaluate our model’s robustness to random perturbations, we adopt the methodology used in previous studies (Haliassos et al. 2021; Zheng et al. 2021), which involves examining four distinct types of degradation: Gaussian blur, Block-wise distortion, Change contrast, and JPEG compression. We apply each of these perturbations at five different levels to assess the model’s robustness under varying degrees of distortion. The video-level AUC results for these unseen perturbations, using the model trained on FF++_c23, are depicted in Fig. 4. Generally, our approach shows higher results than other methods.
4.6 Ablation Study
Effectiveness of our plug-and-play strategy.
Effectiveness of VB and StA.

### v2

@@ Abstract
Three key challenges hinder the development of current deepfake video detection: (1) Temporal features can be complex and diverse: how can we identify general temporal artifacts to enhance model generalization? (2) Spatiotemporal models often lean heavily on one type of artifact and ignore the other: how can we ensure balanced learning from both? (3) Videos are naturally resource-intensive: how can we tackle efficiency without compromising accuracy? This paper attempts to tackle the three challenges jointly. First, inspired by the notable generality of using image-level blending data for image forgery detection, we investigate whether and how video-level blending can be effective in video. We then perform a thorough analysis and identify a previously underexplored temporal forgery artifact: Facial Feature Drift (FFD), which commonly exists across different forgeries. To reproduce FFD, we then propose a novel Video-level Blending data (VB), which is implemented by blending the original image and its warped version frame-by-frame, serving as a hard negative sample to mine more general artifacts. Second, we carefully design a lightweight Spatiotemporal Adapter (StA) to equip a pretrained image model with the ability to capture both spatial and temporal features jointly and efficiently. StA is designed with two-stream 3D-Conv with varying kernel sizes, allowing it to process spatial and temporal features separately. Extensive experiments validate the effectiveness of the proposed methods; and show our approach can generalize well to previously unseen forgery videos, even the latest generation methods.
@@ 1 Introduction
Current deepfake technology enables effortless manipulation of facial identities in video content, bringing many impressive applications in fields such as movie-making and entertainment. However, this technology can also be misused for malicious purposes, including violating personal privacy, spreading misinformation, and eroding trust in digital media. Hence, there is a pressing need to establish a reliable system for detecting deepfake videos.
@@
Most prior deepfake detectors [53, 23, 35, 49] perform well within the same dataset. However, they often fail to generalize well in cross-dataset scenarios where training and testing data distributions differ. Training with synthetic (blending) images appears to be one of the most effective solutions to this problem, as evidenced by [22, 36], which encourages detectors to learn generic representations (e.g., blending boundaries and inconsistencies between inner and outer faces) for deepfake detection. Given the notable success of image-level blending, we ask: what about video-level blending? Can this method also be applied to learn general forgery artifacts in the video domain? (Research-Question-1).
@@
To this end, we perform a comprehensive examination of current realistic forgery videos and discover an inconsistency in the facial organs between forgery frames (as illustrated in Fig. 1). We term this phenomenon as the Facial Feature Drift (FFD). As illustrated in Fig. 1, even when two consecutive frames appear relatively static (possibly without other unrelated effects), subtle unnatural drift and flickering can be observed within facial features (e.g., eyes, nose, etc.). This observation provides reasonable evidence that unnatural forgery artifacts are indeed present in the temporal domain of manipulated videos.
@@
Intuitively, the root cause of FFD can be attributed to the frame-by-frame face-swapping operation employed by existing deepfake algorithms. To understand the underlying reasons, we argue that the primary factor is the randomness during the generation process of fake faces. Deepfake algorithms typically create a fake frame by using DNNs to generate a counterfeit inner face. Subsequently, they perform a blending operation to integrate this generated face back into the original background. Given this context, it is important to note that the inner face between different frames within a single generated video may exhibit variations, as the DNN is not a deterministic function, thereby resulting in inconsistencies during the generation process of fake faces.
@@
To effectively encourage the detection model to capture FFD in the given deepfake, we introduce a novel Video-level Blending data called VB to simulate and reproduce FFD. The overall creation pipeline of VB is illustrated in Fig. 2. Specifically, we generate VB by first applying warping operations separately to facial organs and then performing image-level self-blending to generate one synthesis image. The process is repeatedly conducted for all frames within one clip to create a set of synthesis images (video-level pseudo-fakes). VB can serve as the hard negative samples for mining more general discriminative features for detection. But why do we specifically focus on facial features? This is mainly because regions, such as skin, are low-level areas that can be sensitive to random perturbations (e.g., compression), while facial features typically possess a more robust high-level layout. Also, we can observe more obvious temporal inconsistency within the facial organs in Fig. 1.
@@
Additionally, as suggested by [52, 43], naively training a temporal detector may result in the model only focusing on easier types of artifacts and overlooking more challenging ones (e.g., learning spatial only). So how do we jointly learn temporal and spatial features without heavily relying on one? (Research-Question-2). Also, as videos are naturally more complex than images with multiple frames to be computed, how do we address the efficiency issue? (Research-Question-3). Instead of designing a new video architecture from scratch, we propose an alternative approach: to “convert” a pretrained image model into a video model, enabling the model to learn additional temporal information. Specifically, we meticulously design a lightweight, plug-and-play Spatiotemporal Adapter (StA) that equips an image model (e.g., ViT or CNN) with spatiotemporal reasoning capabilities. Throughout the whole training process, only the added adapters are updated, which significantly reduces the training cost and allows us to leverage the powerful generalization capabilities of SoTA image models (e.g., CLIP).
@@ 2.1 Deepfake Image Detection
Conventional image detectors [31, 35] typically focus on developing optimal CNN architectures (e.g., MesoNet, CapsuleNet). However, these methods often overlook the details present in the frequency domain of fake images, such as compression artifacts. To this end, several works [33, 12] utilize frequency information to improve the performance of detectors. Other notable directions are focusing on some specific representations, i.e., forgery region location [30], 3D decomposition [55], disentanglement [25, 46], erasing technology [41], theoretic information [37], and attentional networks [6, 42]. These works typically utilize designed algorithms for detecting deepfakes from the view of representation learning. Another important direction of image forgery detection is data synthesis (image blending). We will discuss it separately in another subsection.
Methods Blending Mode Introduced Artifacts
Image Level Video Level Face Level Region Level
Face Xray [22], PCL [51], SBI [36], OST [3]
SLADD [2]
SeeABLE [21]
Ours
Table 1: Comparison of our data synthesis method and other similar methods. Ours clearly differs from previous methods in terms of the level at which blending modes are focused (image-level vs. video-level) and the level at which artifacts are introduced (face-level vs. region-level).
Symbol Definition
$\mathbf{I}$ Original image/frame from the original video.
$\mathbf{I}^{\prime}$ Final blended image/frame.
[…]
$\mathbf{I}^{\prime}_{r}$ Blended image/frame about region $r$.
$\mathbf{I}_{\text{warp}}$ Warped version of $\mathbf{I}$ obtained through affine transformations.
$\mathbf{L}_{z}$ Set of landmarks defining region $z,z\in R\bigcup\{R\}$.
$\mathbf{L}^{*}_{R}$ Set of landmarks defining region $R$ after perturbation.
$u_{r}$ Maximum rotation angle for affine transformations.
$u_{s}$ Maximum scaling factor for affine transformations.
$u_{t}$ Maximum translation distance for affine transformations.
$\mathbf{M}_{r}(x,y)$ Mask for region $r$, indicating the blending weight at pixel $(x,y)$.
$\text{dist}(x,y,\mathbf{L}_{r})$ Signed distance function calculating the shortest distance from a pixel $(x,y)$ to the closest landmark in $\mathbf{L}_{r}$.
$\text{fdist}_{r}$ Predefined fall-off distance for the mask of region $r$, controlling the softness of mask edges.
$\alpha_{r}$ Blending weight for region $r$, determining its influence in the final composite image.
$\text{clip}(x,min,max)$ Function that limits $x$ within the range $[min,max]$.
Table 2: Variables and functions definitions for the proposed VB.
@@ 2.2 Deepfake Video Detection
There are parts of works [53] that focus on temporal artifacts. An early work [16] (in 2021) proposes to leverage the temporal inconsistency of the mouth movement. This work utilizes a pre-trained model to learn the natural lip movement as prior on a lip-reading dataset. FTCN [52] explores directly training a fully temporal 3D ConvNets with an attached temporal Transformer. However, detecting without spatial information may harm the generalization capability. STIL [13] considers both spatial and temporal inconsistency and designs a spatiotemporal inconsistency Learning framework for deepfake video detection. RealForensics [17] introduces audio information and leverages self-supervised learning for representation learning. Very recently, wang et al. [43] propose a new learning method that alternatively trains and freezes the spatial and temporal convolutional layers in a 3D CNN (e.g., I3D R50). The result shows the importance of leveraging both spatial and temporal features for detecting deepfakes.
@@ 2.3 Deepfake Detectors Based on Data Synthesis
Image-level Data Synthesis. One effective approach in deepfake detection is to use synthetic data (or blending data/images) for training. This strategy encourages models to learn generic representations for detecting deepfakes, such as blending boundaries and inconsistency between the inner and outer faces [36]. For instance, in the early stages, FWA [23] employs a self-blending strategy by applying image transformations (e.g., down-sampling) to the facial region and then warping it back into the original image. This process is designed to learn the wrapping artifacts during the deepfake generation process. Another noteworthy contribution is Face X-ray [22], which explicitly encourages detectors to learn the blending boundaries of fake images. Furthermore, SLADD [2] introduces an adversarial method to dynamically generate the most challenging blending choices for synthesizing data. Rather than swapping faces between two different identities, a recent art, SBI [36], proposes to swap with the same person’s identity to reach a high-realistic face-swapping. Given the notable generalization performance for detecting image forgery by using synthetic data, it remains an open question whether and how video-level synthetic data can be effective.
@@
Video-level Data Synthesis. Compared to image-level data synthesis (blending), very little works pay attention to designing a suitable video-level blending method. To the best of our knowledge, we identify one notable work [43] that attempts to design video-level data synthesis by using techniques such as dropping/repeating one frame within one clip. However, it is possible that these common operations might not explicitly represent the general forgery pattern. Recall that in the image domain, we use blending technology to synthesize the forgery artifact since the blending artifact commonly exists across different forgeries, generally. So there are two main challenges: (1) What kinds of features generally exist in the temporal domain? (2) How to use blending to simulate it like the image data synthesis? We will show our video-level method in the following section.
@@ 3.3 Spatiotemporal Adapters (StA)
A video can be naturally represented as a stack of image frames. One challenging problem is to capture both spatial and temporal forgery artifacts without heavily relying on one type of artifact. Previous works generally follow two mainstream frameworks: (1) Utilizing 3D CNN (e.g., I3D R50) for capturing spatiotemporal information from forgery videos [52, 43]; (2) Utilizing two-stream architecture to learn the spatial and temporal features in two separate branches [29, 53]. These two paradigms typically involve fully fine-tuning for both spatial and temporal modules. However, training all layers of both spatial and temporal modules will inevitably introduce a large number of learnable parameters, making the optimization and training procedure challenging. On the other hand, the image model (especially ViT-based image SoTA) is developing rapidly. It is well-known that these image SoTAs have achieved notable performance in several downstream computer vision tasks (e.g., multi-modality [11]). Therefore, can we leverage the power of these SoTA image models to effectively and efficiently finetune the face forgery detection task? The key challenge is to perform the temporal forgery feature modeling in a given image model.
@@
4 Results
4.1 Setup
Table 3: Protocol-1: Cross-dataset evaluations with image detectors. All detectors are trained on FF++_c23 [35] and evaluated on other datasets. The best results are highlighted in bold and the second is underlined. * indicates the results are directly cited from their original papers, otherwise the results are cited by [47, 4, 27].
@@ 4.1 Setup
F3Net [33] ECCV 2020 0.789 0.844 0.718 0.749 0.730 0.728 0.649 0.743
SPSL [26] CVPR 2021 0.799 0.871 0.724 0.770 0.723 0.702 0.794 0.769
SRM [28] CVPR 2021 0.840 0.885 0.695 0.728 0.722 0.702 0.794 0.767
CORE [32] CVPRW 2022 0.809 0.882 0.721 0.720 0.765 0.724 0.710 0.762
RECCE [1] CVPR 2022 0.823 0.891 0.696 0.734 0.784 0.756 0.711 0.779
DCL [38] AAAI 2022 0.882 0.921 0.750 0.769 0.975 0.776 0.863 0.851
SLADD* [2] CVPR 2022 0.797 - 0.772 - - - - -
SBI [36] CVPR 2022 0.886 0.827 0.717 0.848 0.899 0.703 0.866 0.821
UIA-ViT* [56] ECCV 2022 0.889 - - 0.795 - - - -
UCF [46] ICCV 2023 0.837 0.867 0.742 0.770 0.808 0.774 0.697 0.785
SeeABLE* [46] ICCV 2023 0.873 - 0.759 0.863 - - - -
IID* [19] CVPR 2023 0.838 0.939 - 0.812 - - - -
CFM [27] TIFS 2023 0.897 0.952 0.706 0.802 0.976 0.823 0.831 0.856
LSDA [45] CVPR 2024 0.898 0.956 0.735 0.812 0.892 0.756 0.879 0.847
@@
Table 4: Protocol-2: Cross-manipulation evaluations within FF++ domain. All models are trained on FF++ (c23) and tested on other forgeries. $\dagger$ indicates the results are obtained by using the official pre-trained model for evaluation, otherwise the results are cited by [47, 4].
@@
RECCE [1] 0.898 0.832 0.925 0.683 0.848 0.949 0.848 0.768 0.844
SBI [36] 0.724 0.891 0.952 0.750 0.594 0.803 0.712 0.701 0.766
UCF [46] 0.831 0.827 0.950 0.731 0.862 0.937 0.809 0.647 0.824
[…]
@@
LipForensics [16] CVPR 2021 0.824 0.735 0.780
FTCN [52] ICCV 2021 0.869 0.740 0.805
HCIL [14] ECCV 2022 0.790 0.692 0.741
RealForensics [17] CVPR 2022 0.857 0.759 0.808
LTTD [15] NeurIPS 2022 0.893 0.804 0.849
AltFreezing [43] CVPR 2023 0.895 - -
TALL-Swin [44] ICCV 2023 0.908 0.768 0.838
StyleDFD [5] CVPR 2024 0.890 - -
NACO [50] ECCV 2024 0.895 0.767 0.831
@@
I3D-RGB [16] 25M 0.625 0.655 0.640
3D R50 [16] 46M 0.824 0.735 0.780
VTN [13] 46M 0.734 0.691 0.713
VidTR [52] 93M 0.869 0.740 0.805
ViViT [43] 310M 0.895 - -
ISTVT [40] - 0.606 0.551 0.579
ResNet-34 [18] 21M $\times$ 0.742 0.626 0.684
@@
Implementation Details. We use the CLIP ViT-l14 [34] as the default backbone. We also explore other architectures for experiments, as shown in Tab. 6. All preprocessing and training codebases are aligned to the deepfake benchmark [47]. Specifically, we uniformly sample 32 frames from each video during training and randomly select 8 consecutive frames as a clip for each iteration. We also employ several widely used data augmentations, such as CutOut, and ImageCompression.
Evaluation Protocols and Dataset. We adopt three standard protocols to evaluate our models: Protocol-1: cross-dataset evaluation with image methods, Protocol-2: cross-manipulation evaluation on the FF++ data domain, and Protocol-3: cross-dataset evaluation with video methods. For Protocol-1 and Protocol-3, we evaluate the models by training them on FaceForensics++ (FF++) [35] and then testing them on seven other deepfake detection datasets. Note that FF++ has three different compression versions and we adopt the c23 version for training all methods in our experiments, following most existing works [45]. The evaluation datasets include Celeb-DF-v2 (CDF-v2) [24], DeepfakeDetection (DFD) [7], Deepfake Detection Challenge (DFDC) [9], the preview version of DFDC (DFDCP) [8], DeeperForensics (DFo) [20], WildDeepfake (WDF) [57], and FFIW [54]. For Protocol-2, we evaluate the models using the latest deepfake dataset DF40 [48], which contains diverse forgery data generated within the FF++ domain. By doing this, we can ensure that although the evaluation involves different forgeries, the data domain remains the same.
4.2 Generalization Performance Evaluation
Forgery methods in the real world can be highly diverse. Evaluating our model’s performance in diverse scenarios is needed. We conduct both cross-dataset evaluations (Protocol-1 and Protocol-3) and cross-manipulation tasks, allowing us to determine whether the model can handle both previously unseen data domains and unseen fake methods. Results in Fig. 3, Fig. 4, and Fig. 5 show that our method generally outperforms other models on average, particularly on e4s, Inswap, and SimSwap methods (see Tab. 4). This demonstrates that our method learns more general features for detection, even for the latest techniques.
Figure 4: GradCAM for demonstrating the attention regions of S- and T-Adapters. We create a “boring” video by repeating a single image into a video sequence. We show that the T-adapter can capture reasonable temporal-related motion like mouth movement, while the S-adapter’s outputs remain constant due to the same input of the boring video.
Figure 5: Comparison of different video-level blending strategies. We consider other two possible solutions to simulate the FFD: (1) CBI, which represents the face-hull blending of two frames with the same video clip; (2) PFIG [39], which represents the facial-region blending of the different videos. Our simulation method VB shows the most similar result to the original FFD in the fake data. Best viewed in color.